### v1

# LayLens: Improving Deepfake Understanding through Simplified Explanations

Abstract.
This demonstration paper presents LayLens, a tool aimed to make deepfake understanding easier for users of all educational backgrounds. While prior works often rely on outputs containing technical jargon, LayLens bridges the gap between model reasoning and human understanding through a three-stage pipeline: (1) explainable deepfake detection using a state-of-the-art forgery localization model, (2) natural language simplification of technical explanations using a vision-language model, and (3) visual reconstruction of a plausible original image via guided image editing. The interface presents both technical and layperson-friendly explanations in addition to a side-by-side comparison of the uploaded and reconstructed images. A user study with 15 participants shows that simplified explanations significantly improve clarity and reduce cognitive load, with most users expressing increased confidence in identifying deepfakes. LayLens offers a step toward transparent, trustworthy, and user-centric deepfake forensics.
Deepfake Detection, Explanation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; July 2025; Melbourne, Australiabooktitle: ICMI ’25: INTERNATIONAL CONFERENCE ON MULTIMODAL INTERACTION, October 13–17, 2025, Canberra, Australiasubmissionid: 1003ccs: Human-centered computing Graphical user interfaces
1. Introduction
The advancement in generative AI technology has led to a proliferation of AI-generated and manipulated content (commonly referred to as deepfakes), raising significant challenges for media authenticity, public trust and digital safety. While some of the state-of-the-art deepfake detection tools provide textual explanations of why an image may be a fake, their outputs are often opaque and overly technical for non-expert users. To address this critical gap, we present LayLens: an intuitive, web-based interface that allows users to upload an image and receive highly simplified, visually guided explanations of why the image may be fake, along with a plausible reconstruction of what the authentic image might have looked like. Our goal is to bridge the gap between high-performance deepfake detection and public interpretability by offering explanations that are not only accurate but also immediately understandable to general audiences, including educators, journalists, content moderators and everyday users.
2. Related Work
Traditional deepfake detection models, such as  (Zhou et al., 2018; Matern et al., 2019; Chugh et al., 2020) primarily focused on obtaining high accuracy in detecting manipulated inputs but lacked intuitive interpretability for lay users. In earlier years, explainability in deepfake detection was introduced through saliency-based techniques like LRP (Bach et al., 2015), Grad-CAM (Selvaraju et al., 2017), LIME (Ribeiro et al., 2016) and SHAP (Lundberg and Lee, 2017), which produce saliency maps highlighting image regions most influential to the classifier’s output. These approaches are model-agnostic and were adopted to identify the important input regions in deepfake detection models in works such as  (Pino et al., 2021; Malolan et al., 2020). Later works such as  (Soltandoost et al., 2025) propose identifying embeddings corresponding to local facial regions in images, and use them to produce interpretable classification outputs. Recently, Vision-Language model based approaches, such as FakeShield  (Xu et al., 2025) and SIDA  (Huang et al., 2025) have become popular, which provide textual rationale along with localization of the tampered region. Despite these advancements, current explainable detection tools exhibit gaps that limit their usefulness for non-experts. A common issue is that the explanations are too low-level or technical. There is an evident gap in the literature for a deepfake detection interface that provides simplified yet informative visual and textual explanations accessible to the general public.
3. Interface Design
Figure 1. Comparison View: The user-uploaded fake image is analyzed by the system to localize regions suspected of manipulation (left). Based on these findings, a plausible reconstruction of the original, unaltered image is generated (right). This side-by-side view facilitates intuitive visual understanding of the manipulated content.
[…]
Figure 2. Complex Explanation: This view provides a detailed breakdown of several technical factors such as lighting inconsistencies, resolution artifacts, perspective anomalies, and shadow discrepancies that inform the model’s decision in identifying potential manipulations within the image.
[…]
Figure 3. Simplified Explanation: This view translates the technical explanation into concise, region-level descriptions tailored for non-expert users. By reducing cognitive load and using accessible language, it enhances interpretability and user engagement.
[…]
LayLens is designed as an interactive, end-to-end system that not only detects deepfake images but also presents explanations in a form accessible to both technical and non-technical users. The system integrates state-of-the-art methods in deepfake detection, localization and explanation, along with generative image editing, and wraps them in a user-friendly interface that encourages interpretability and engagement. After the user uploads an image to the interface, the following components are triggered:
3.0.1. Comparison View (Figure 1):
First, the users are presented with a side-by-side visual display. The left panel shows the original uploaded image, overlaid with a softly pulsating mask which highlights regions suspected to have been manipulated. The right panel shows a reconstructed version of the image, generated based on the AI’s understanding of what the non-tampered image could have looked like. A slider enables intuitive, pixel-level comparison between the two. While the mask is obtained through Fakeshield (Xu et al., 2025), Step1X-Edit (Liu et al., 2025) creates the imagined version of the original image, by taking the user-uploaded image along with the edit instruction (as obtained in Section 3.0.2(2) below) as inputs. We also tried using ICEdit (Zhang et al., 2025) here, but found Step1X-Edit’s editing performance to be better (based upon manual observation).
3.0.2. Explanation Card:
Below the comparison view, users can access a flip-style explanation card. This card presents the system’s reasoning in two tiers:
  1. (1)
    Complex Explanation (Figure 2): Offers detailed reasoning behind the detection decision, including references to lighting inconsistencies, perspective errors, shadow artifacts, resolution discrepancies and physically implausible structures. This is obtained from the output of the Fakeshield (Xu et al., 2025) model. Here, we also experimented with SIDA (Huang et al., 2025), but found Fakeshield’s explanations to be more intuitive and accurate through manual observation.
  2. (2)
    Simplified Explanation (Figure 3): Translates the technical rationale into natural language, tailored for non-expert users. Each identified region (e.g., “Central Teddy Bear”) is paired with a relevant emoji and an accessible explanation. Specifically, we employ Qwen-VL (Bai et al., 2025), a foundation Vision Language Model. It is prompted with both the image and FakeShield’s technical explanation and tasked with generating a structured, human-readable JSON output. For each manipulated region, Qwen-VL outputs:
    • A simplified explanation of why the region appears fake,
    • An associated emoji to visually cue the user,
    • A concise edit instruction describing how to correct or restore the manipulated region.
Thus, LayLens offers a complete workflow that adapts to users’ preferred level of technical detail while always grounding its decisions in visual evidence. It enhances the accessibility, transparency, and interpretability of deepfake detection for a broad spectrum of users, from forensic analysts to everyday citizens. This design lets different users choose their preferred explanation detail level, while always seeing a visual demonstration of the AI’s reasoning.
4. User Survey
Figure 4. Distribution of User survey likert scale responses for various questions.
[…]
To evaluate the effectiveness of LayLens in providing accessible explanations for deepfake images, we conducted a user study with 15 participants, of whom 11 were familiar with the concept of deepfakes, and 3 had prior experience with deepfake detection tools. Each participant interacted with the system by analyzing 10 AI-manipulated images. Overall, users preferred the simplified explanations over the complex ones in 65.3% of the cases. Notably, in 81.3% of comparisons, participants reported that the simplified explanations reduced their cognitive load in understanding why an image might be a deepfake. The side-by-side visualization, featuring the uploaded (potentially fake) image alongside a plausibly reconstructed original, was considered helpful in enhancing understanding in 69.3% of the instances. Furthermore, 80% of participants indicated that the experience improved their confidence in detecting deepfakes in the future, and 93.3% expressed interest in using such a tool for identifying manipulated media going forward. The distribution of responses across all questions, measured using a 5-point Likert scale, is illustrated in Figure 4. Additionally, we performed a Wilcoxon signed-rank test to assess perceived changes in Ease of Understanding, Clarity and Accuracy when switching from complex to simplified explanations. The resulting p-values were $063.25e-$ 06, $0.01$ and $0.30$, respectively. These results indicate statistically significant improvements (p $\leq 0.05$) in both Ease of Understanding and Clarity when simplified explanations are presented.
5. Conclusion
In this work, we presented LayLens, a system to make deepfake detection more accessible, interpretable and engaging for users with varying levels of technical expertise. By integrating detection, natural language simplification, and generative reconstruction, LayLens enables both technical and non-expert users to understand why an image may be fake and what the original might have looked like. Our user study shows that simplified, visually grounded explanations reduce cognitive load and enhance user confidence. LayLens demonstrates that deepfake detection can be made both accurate and accessible, paving the way for more transparent and trustworthy media forensics. In future, we will add performance comparisons of this tool using proprietary AI tools such as ChatGPT (OpenAI, 2024) and Gemini (Anil and et al., 2023); and work towards creating an open-source, efficient VLM to directly obtain simplified explanations.

### v2

@@
 LayLens offers a step toward transparent, trustworthy, and user-centric deepfake forensics.
-Deepfake Detection, Explanation
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; July 2025; Melbourne, Australiabooktitle: ICMI ’25: INTERNATIONAL CONFERENCE ON MULTIMODAL INTERACTION, October 13–17, 2025, Canberra, Australiasubmissionid: 1003ccs: Human-centered computing Graphical user interfaces
+Deepfake Detection; Explanation
+$c$s: Human-centered computing Graphical user interfacesccs: Computing methodologies Biometricsccs: Human-centered computing Accessibility systems and toolsccs: Social and professional topics Assistive technologies
 1. Introduction
@@
 Traditional deepfake detection models, such as  (Zhou et al., 2018; Matern et al., 2019; Chugh et al., 2020) primarily focused on obtaining high accuracy in detecting manipulated inputs but lacked intuitive interpretability for lay users.
-In earlier years, explainability in deepfake detection was introduced through saliency-based techniques like LRP (Bach et al., 2015), Grad-CAM (Selvaraju et al., 2017), LIME (Ribeiro et al., 2016) and SHAP (Lundberg and Lee, 2017), which produce saliency maps highlighting image regions most influential to the classifier’s output.
+In earlier years, explainability in deepfake detection was introduced through saliency-based techniques like LRP (Bach et al., 2015), Grad-CAM (Selvaraju et al., 2017), LIME (Ribeiro et al., 2016) and SHAP (Lundberg et al., 2017), which produce saliency maps highlighting image regions most influential to the classifier’s output.
 These approaches are model-agnostic and were adopted to identify the important input regions in deepfake detection models in works such as  (Pino et al., 2021; Malolan et al., 2020).
@@
 […]
+Figure 5. Distribution of User survey likert scale responses for various questions.
+[…]
 LayLens is designed as an interactive, end-to-end system that not only detects deepfake images but also presents explanations in a form accessible to both technical and non-technical users.
 The system integrates state-of-the-art methods in deepfake detection, localization and explanation, along with generative image editing, and wraps them in a user-friendly interface that encourages interpretability and engagement.
+(Figure 1) gives an overview of this three-stage pipeline, showing how LayLens moves from deepfake detection, through explanation simplification, to visual reconstruction.
 After the user uploads an image to the interface, the following components are triggered:
@@
   2. (2)
-    Simplified Explanation (Figure 3): Translates the technical rationale into natural language, tailored for non-expert users.
-Each identified region (e.g., “Central Teddy Bear”) is paired with a relevant emoji and an accessible explanation.
-Specifically, we employ Qwen-VL (Bai et al., 2025), a foundation Vision Language Model.
+    Simplified Explanation (Figure 4): Performs automatic text simplification on complex explanation.
+Each identified region (e.g., “Central Teddy Bear”) is paired with a relevant emoji and a simplified explanation, generated by a VLM (Bai et al., 2025).
 It is prompted with both the image and FakeShield’s technical explanation and tasked with generating a structured, human-readable JSON output.
-For each manipulated region, Qwen-VL outputs:
+For each manipulated region, the VLM outputs:
     • A simplified explanation of why the region appears fake,
@@
     • A concise edit instruction describing how to correct or restore the manipulated region.
-Thus, LayLens offers a complete workflow that adapts to users’ preferred level of technical detail while always grounding its decisions in visual evidence.
+Thus, LayLens offers a complete workflow that lets the users choose their preferred level of technical detail while always grounding its decisions in visual evidence.
 It enhances the accessibility, transparency, and interpretability of deepfake detection for a broad spectrum of users, from forensic analysts to everyday citizens.
@@
 4. User Survey
-Figure 4. Distribution of User survey likert scale responses for various questions.
-[…]
 To evaluate the effectiveness of LayLens in providing accessible explanations for deepfake images, we conducted a user study with 15 participants, of whom 11 were familiar with the concept of deepfakes, and 3 had prior experience with deepfake detection tools.
@@
 Additionally, we performed a Wilcoxon signed-rank test to assess perceived changes in Ease of Understanding, Clarity and Accuracy when switching from complex to simplified explanations.
-The resulting p-values were $063.25e-$ 06, $0.01$ and $0.30$, respectively.
+The resulting p-values were 3.25e063.25e-06, $0.01$ and $0.30$, respectively.
 These results indicate statistically significant improvements (p $\leq 0.05$) in both Ease of Understanding and Clarity when simplified explanations are presented.
@@
 LayLens demonstrates that deepfake detection can be made both accurate and accessible, paving the way for more transparent and trustworthy media forensics.
-In future, we will add performance comparisons of this tool using proprietary AI tools such as ChatGPT (OpenAI, 2024) and Gemini (Anil and et al., 2023); and work towards creating an open-source, efficient VLM to directly obtain simplified explanations.
+As a future work, we will develop an open-source, efficient and explainable model for large-scale audio-visual deepfakes (Cai et al., 2025) and further, extend it to incorporate multi-lingual, code-switched videos (Kuckreja et al., 2025).